OpenStudy (anonymous):

Solve: x2 - 10x - 24 = 0.

OpenStudy (anonymous):

x=12,-2

Parth (parthkohli):

=> x2 - 12x + 2x - 24 = 0 => x(x - 12) + 2(x - 12) = 0 => (x + 2)(x - 12) = 0 So, 2 and -12 are answers.

Parth (parthkohli):

I mean -2 and 12

OpenStudy (anonymous):

Thanks yall ;D